OpenStudy (anonymous):

Identify the property shown. 6 • w = w • 6 A. transitive B. associative C. commutative D. distributive

OpenStudy (anonymous):

@IMxLORDxGOTEN

OpenStudy (anonymous):

@bohotness

OpenStudy (anonymous):

@e.mccormick

OpenStudy (bohotness):

which one do you think is your answer?

OpenStudy (anonymous):

i think its A.

OpenStudy (bohotness):

acttiy its c or b love

OpenStudy (anonymous):

oh....

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

Do you know the different properties?

OpenStudy (kendricklamar2014):

The answer is Commutative Property. Commutative Property Rule: a + b = b + a

OpenStudy (anonymous):

thanks @KendrickLamar2014

OpenStudy (kendricklamar2014):

If you don't know the definitions of all these terms, here is the definition's and examples: http://www.purplemath.com/modules/numbprop.htm

OpenStudy (kendricklamar2014):

No Problem

OpenStudy (anonymous):

wanna see what i got out of 3??

OpenStudy (anonymous):

100% Assessment Summary Score: 3/3 points Status: Mastered Date Taken: Mon, Mar. 30, 20152:04 PM PDT Retest Information If you retake this assessment, only your new score — and not your previous scores — will count. Review AssessmentRetake Assessment Continue with lesson

TheSmartOne (thesmartone):

The Transitive Property of Equality is If a = b and b = c, then a = c. Distributive property is a(b + c) = ab + ac Associative property is a + (b + c) = (a + b) + c commutative property is ab = ba And use that to get your answer, which you have already gotten :P